OpenStudy (anonymous):

if a^(-b)=1then 1/a^b = 1 a^b=1 thus a^b = a^(-b) or, b = -b what is wrong

OpenStudy (lgbasallote):

which do you think is wrong?

OpenStudy (anonymous):

b cannot be equal to -b. so what I did must have some type of error.

OpenStudy (lgbasallote):

why cant it? a^b = 1 a^(-b) = 1 therefore a^b = a^(-b)

OpenStudy (lgbasallote):

according to a property \[\text{ if} \; a^m = a^n \; \text{then} \; m = n\] are you familiar with that?

OpenStudy (anonymous):

yes. but if b = 3 then how can 3=-3 be true

OpenStudy (anonymous):

oh. i think i found the answer.

OpenStudy (anonymous):

nothing is wrong. but what i thought was mistake b cannot be any other number than 0.so 0=-0. sorry for the wrong question.

OpenStudy (lgbasallote):

haha yup

OpenStudy (lgbasallote):

i wanted to tell you that but i figured finding out on your own was more fun :S